OpenStudy (anonymous):

if the joint probability density of X and Y is given by f(x,y)= .25(2x+y) for 0

OpenStudy (anonymous):

its a double integral so take the integral with respect to x with the limits of x and then take that answer and do a second integral with respect to y and use those limits

OpenStudy (anonymous):

since I'm only looking for the marginal density of X the book says g(x)=\[\int\limits_{?}^{?}\]f(x,y)dy for \[-\infty\]<x<\[\infty\]....so its not a double integral in this case...but when I did the problem i used the 0<x<1 but my answer was incorrect, so I'm thinking that maybe i should've used the 0<y<2...any suggestions?

OpenStudy (anonymous):

ok wow I didn't know it was going to start a new paragraph everytime I used the equation editor

OpenStudy (anonymous):

i would try that since your taking integral with respect to y

OpenStudy (anonymous):

It is a double integral first evaluate the integral of y, then evaluate the integral of x\[\int\limits_{0}^{1}\int\limits_{0}^{2}(1/2x+1/4y)dydx\]